# How Transliterations Improve Crosslingual Alignment

### v1

@@ Abstract
Recent studies have shown that post-aligning multilingual pretrained language models (mPLMs) using alignment objectives on both original and transliterated data can improve crosslingual alignment. This improvement further leads to better crosslingual transfer performance. However, it remains unclear how and why a better crosslingual alignment is achieved, as this technique only involves transliterations, and does not use any parallel data. This paper attempts to explicitly evaluate the crosslingual alignment and identify the key elements in transliteration-based approaches that contribute to better performance. For this, we train multiple models under varying setups for two pairs of related languages: (1) Polish and Ukrainian and (2) Hindi and Urdu. To assess alignment, we define four types of similarities based on sentence representations. Our experiments show that adding transliterations alone improves the overall similarities, even for random sentence pairs. With the help of auxiliary alignment objectives, especially the contrastive objective, the model learns to distinguish matched from random pairs, leading to better alignments. However, we also show that better alignment does not always yield better downstream performance, suggesting that further research is needed to clarify the connection between alignment and performance.
@@ 1 Introduction
The training of highly multilingual language models has to cope with the diversity of scripts (e.g. more than 30 in Glot500-m (ImaniGooghari et al., 2023)), which tends to reduce the effectiveness of crosslingual transfer (Dhamecha et al., 2021; Purkayastha et al., 2023). A few recent works explore the possibility of using a common script to represent all languages (Purkayastha et al., 2023; Moosa et al., 2023) through transliteration, a process of converting the text of a language from one script to another script (Wellisch et al., 1978). The intuition is that a common script can help the model to learn more knowledge through lexical overlap since common vocabularies have been shown to contribute to better crosslinguality (Pires et al., 2019; Amrhein and Sennrich, 2020). However, this script normalization step yields models that only support one script. This also hinders efficiency, as texts have to be transliterated into the common script before fed to the model. In addition, transliteration ambiguity (words with different meanings having the same transliteration) can be a potential problem (Liu et al., 2024c).
@@
Instead of relying solely on common-script transliterations, recent works also use transliterations as an auxiliary input to improve crosslingual alignment without expanding the vocabulary (Liu et al., 2024b; Xhelili et al., 2024). These approaches combine sentences in their original script alongside with their transliteration as paired inputs for sentence- or token-level alignment objectives. Surprisingly, even without parallel data, these methods show remarkable improvement in crosslingual transfer between languages with different scripts. However, the concept of crosslingual alignment is often vaguely defined in these studies. Moreover, it remains unclear why incorporating transliterated data and auxiliary alignment objectives contributes to better crosslingual alignment, which relates to the similarity between translation pairs.
@@
To this end, this work presents – to the best of our knowledge – the first attempt to explain why using transliteration as data augmentation and incorporating transliteration-based alignment objectives, such as transliteration contrastive learning (TCM) (Liu et al., 2024b) and transliteration language modeling (TLM) (Xhelili et al., 2024), can improve the crosslingual alignment. We first discuss definitions of crosslingual alignment (Roy et al., 2020; Hämmerl et al., 2024) and establish a clear connection between crosslingual alignment and sentence retrieval performance. We then conduct a case-study on two related language pairs crucially using different scripts: (a) Polish-Ukrainian and (b) Hindi-Urdu, to explore how and why the transliteration-augmented approach improves crosslingual alignment. Specifically, we aim to answer the following questions: (1) Does adding transliterated data alone improve alignment? (2) How do auxiliary objectives contribute to better alignment? (3) How does alignment vary when the target language is in the original script, the Latin script, or when both the source and target languages are in the Latin script? (4) Does better alignment always lead to better downstream crosslingual performance?
@@
To answer these questions, we define four types of similarities and conduct a thorough analysis of how these similarities vary across multiple model variants. Our key experimental findings can be summarized as follows:
@@
(i) Adding transliterations alone does not improve alignment but does enhance all types of similarities. This occurs as the similarity between randomly paired sentences also increases. However, effective alignment requires distinguishing matched pairs from random pairs. (ii) With auxiliary alignment objectives, transliterations serve as an intermediary. Language 1 in its original script is aligned with the transliterations of Language 1; Language 2 in its original script is aligned with the transliterations of Language 2; transliterations of Language 1 are aligned with those of Language 2. Through this process, the alignment between Language 1 and Language 2 (both in their original scripts) is improved. (iii) Although crosslingual alignment is generally considered crucial for enhancing downstream crosslingual performance, our results indicate that better alignment does not always yield better downstream performance. This finding suggests that further research is needed to clarify the connection between alignment and performance (Hua et al., 2024).
@@ 2.1 Multilingual Language Models
Models pretrained on a wide range of languages using self-supervised objectives, such as masked language modeling (MLM) (Devlin et al., 2019) or causal language modeling (Radford et al., 2019), are referred to as mPLMs. With respect to their use of the Transformer (Vaswani et al., 2017) architecture, these models can be categorized into encoder-only (Devlin et al., 2019; Conneau et al., 2020; Liang et al., 2023), encoder-decoder (Liu et al., 2020; Fan et al., 2021; Xue et al., 2021), and decoder-only models (Lin et al., 2022; Shliazhko et al., 2022; Scao et al., 2022). With the recent scale-up in both model and data size, decoder-only models, also known as large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023), were able to achieve impressive performance in various generation tasks across high- and medium-resource languages (Zhao et al., 2024a; Üstün et al., 2024; Zhao et al., 2024b). Parallel efforts have produced encoder-only models with very large language coverage, improving the situation for many for low-resource languages (Ogueji et al., 2021; Alabi et al., 2022; ImaniGooghari et al., 2023; Liu et al., 2024a). These encoder-only models excel in multiple tasks in a zero-shot crosslingual context (Huang et al., 2019; Artetxe and Schwenk, 2019; Hu et al., 2020).
@@ 2.2 Training with Alignment Objectives
Most mPLMs trained without any additional crosslingual signals already show good performance across languages, possibly due to factors such as lexical overlap (Pires et al., 2019), shared position and special token embeddings (Dufter and Schütze, 2020) and even language imbalance (Schäfer et al., 2024). A recent review of factors contributing to crosslingual transfert is in (Philippy et al., 2023). To further improve the crosslingual alignment of mPLMs, many methods additionally leverage crosslingual signals during or after pretraining. These methods can rely on bilingual dictionaries (Cao et al., 2020; Wu and Dredze, 2020; Chi et al., 2021b; Efimov et al., 2023), parallel data (Reimers and Gurevych, 2020; Pan et al., 2021; Wang et al., 2022b), or a combination of both (Wei et al., 2021; Hu et al., 2021) to facilitate crosslingual alignment. This set of methods aims to increase the similarity between paired instances (words or sentences) and sometimes also to reduce the similarity between unrelated data, via contrastive learning objectives (Chopra et al., 2005; Gao et al., 2021; Chi et al., 2021a). Another group of methods focuses on reformulating the training data, with the expectation of implicitly improving alignment through techniques such as artificial code-switching generation (Chaudhary et al., 2020; Wang et al., 2022a; Reid and Artetxe, 2022) or using a translation language modeling objective (Conneau and Lample, 2019).
@@ 2.3 Transliteration in Language Modeling
Transliteration converts the text of a language from one script to another (Wellisch et al., 1978). Since this process does not translate meanings but rather represents the original sounds as closely as possible in the target script, transliteration can be performed efficiently and accurately using a rule-based system, such as Uroman (Hermjakob et al., 2018). Recent studies have shown that better language models can be trained using data transliterated into a common script, due to improved lexical overlap (Amrhein and Sennrich, 2020; Dhamecha et al., 2021; Muller et al., 2021; Purkayastha et al., 2023; Moosa et al., 2023). To further address the script barrier and prevent the model from only supporting only one script, another line of work uses transliterations as auxiliary input to construct paired data for post-pretraining fine-tuning (Liu et al., 2024b; Xhelili et al., 2024), leading to better crosslingual performance between languages written in different scripts. To further address the script barrier and prevent models from supporting only one script, another line of research uses transliterations as auxiliary input to create paired data for post-pretraining fine-tuning (Liu et al., 2024b; Xhelili et al., 2024), resulting in better crosslingual performance between languages written in different scripts. However, it remains unclear why these approaches achieve better alignment using only automatic transliterations as auxiliary inputs.
@@ 3 Prelimilary: Crosslingual Alignment
Definition. Crosslingual alignment refers to the degree of similarity among representations of similar meanings across languages, which can be further classified into weak alignment and strong alignment (Roy et al., 2020).111Strong alignment requires even a greater distance of dissimilar meanings within a language, which is usually hard in multilingual NLP. Therefore it is not considered in this paper. Following the definition of weak alignment of Hämmerl et al. (2024), similar meanings (across languages) should have more similar representations than dissimilar meanings. Formally, let be a pair of representations of two units (either tokens or sentences) with similar meanings in language and respectively, weak alignment is defined as follows:
where sim is a similarity measure, e.g., cosine similarity. Weak alignment requires that the representational similarity between a unit in and its (roughly) equivalence in should be higher than the similarities between this unit and any other ones in . It is important to note that this notion of alignment emphasizes the relative magnitude of similarity rather than the absolute magnitude. The similarity between and does not have to be very large to induce compliant alignments. Some models, though assigning similar representations to similar units, also make less related or even unrelated units similar, therefore possibly resulting in . This naturally induces suboptimal alignment due to the failure to differentiate between similar meanings and dissimilar meanings across languages.
Evaluations. Crosslingual alignment can be directly evaluated by sentence retrieval, where a model retrieves the most relevant or similar sentence in language given a query sentence in language . The performance of sentence retrieval can be measured by computing the top- accuracy for a given parallel corpus, using sentences from one language as the queries, and retrieving matched sentences in the other language. Additionally, crosslingual alignment is believed to be able to be evaluated indirectly by other downstream tasks that require zero-shot crosslingual transfer ability (Huang et al., 2019; Artetxe and Schwenk, 2019). The intuition is that models with good crosslingual alignment are often expected to present good transfer performance as the representations of similar meanings are similar across languages. However, we show that better crosslingual alignment does not always induce better downstream crosslingual performance in §5.4, which is also consistent with some previous findings (Wu and Dredze, 2020; Gaschi et al., 2023).
@@ 4.1 Languages
Polish-Ukrainian pair. Polish (pol) and Ukrainian (ukr) are Slavic languages, belonging to the West and East Slavic branch respectively. Polish and Ukrainian have historically influenced each other, contributing to shared vocabulary and linguistic features. Polish uses Latin (Latn) script while Ukrainian uses Cyrillic (Cyrl) script.
Hindi-Urdu pair. Hindi (hin) and Urdu (urd) both belong to the Indo-Aryan branch of the Indo-European language family, spoken in the Indian subcontinent. They are mostly mutually intelligible languages that historically can be viewed as two standardized dialects of Hindustani, and therefore they share extensive common vocabularies. A major difference is that Hindi uses Devanagari (Deva) script while Urdu uses Arabic (Arab) script.
An important difference between these two language pairs is that transliteration only changes the script of one language (ukr) for pol-ukr, whereas it changes the script of both urd and hin for hin-urd.
@@ 4.2 Training Data
Original data. We use the data from Glot500-c (ImaniGooghari et al., 2023) for each language of interest. For the pol-ukr pair, there are around 7M sentences for ukr_Cyrl and around 19M sentences for pol_Latn sentences. For the hin-urd pair, there are around 7M sentences for hin_Deva and 6M sentences for urd_Arab. We concatenate all data together for each language pair and refer to the final data in their original script as Data for pol-ukr and Data for hin-urd.
Transliterated data. We use Uroman (Hermjakob et al., 2018) to transliterate both Data and Data to Latin. We refer to the resulted Latin script data as Data and Data for pol-ukr and hin-urd pair respectively. It is important to note that the original and transliterated data are in one-to-one correspondence. For example, th line in Data is the transliteration of the th line in Data.
4.3 Training Objectives
Masked Language Modeling (MLM). This is the primary learning objective we use to train our model variants. This objective improves the general language modeling ability by masking certain tokens in the input sentences and learning to predict them. Following Devlin et al. (2019), we randomly replace 15% tokens in the input sentences with a special token: [mask] and use a language modeling head to reconstruct the original tokens from the final contextualized embeddings.
Transliteration Contrastive Modeling (TCM). This contrastive objective is proposed by Liu et al. (2024b). It improves the similarity between pairs of sentence-level representations composed of one sentence in its original script and the corresponding Latin transliteration. Following Liu et al. (2024b), we obtain the representations by mean pooling the output of the 8th Transformer layer and calculate the loss batch-wise: the positive samples are the paired sentences within a batch; the negative samples are any combinations of two sentences that are not paired within a batch.
Transliteration Language Modeling (TLM). This objective, proposed by Xhelili et al. (2024), is similar to the translation language modeling of Conneau and Lample (2019), where we use transliterations, instead of translations, to build sentence pairs in the objective. Following (Xhelili et al., 2024), we concatenate a sentence and its transliteration and perform MLM on the combined text. To predict a token masked in the original sentence, the model can either attend to tokens in the original script or their transliterations, and vice versa.
@@ 4.3 Training Objectives
4.4 Models
We train a SentencePiece Unigram tokenizer (Kudo, 2018; Kudo and Richardson, 2018) on Data and Data for each language pair, respectively. We set the size of vocabulary to 30K for each pair. The tokenizers are not adapted to the transliterated data, i.e., Data and Data, in order to replicate the settings used by Liu et al. (2024b) and Xhelili et al. (2024) as they achieve surprisingly good performance without any tokenizer adaptation. As shown in Table 1, lexical overlap increases drastically for transliterated data even without learning subwords from it. We then train five model variants from scratch for each language pair to thoroughly explore the effect of each component of the transliteration-augmented training. We introduce the 5 model variants as follows (training details are reported in §A).
Model-1. These models are trained on either Data or Data only with MLM.

### v2

@@ Abstract
Recent studies have shown that post-aligning multilingual pretrained language models (mPLMs) using alignment objectives on both original and transliterated data can improve crosslingual alignment. This improvement further leads to better crosslingual transfer performance. However, it remains unclear how and why a better crosslingual alignment is achieved, as this technique only involves transliterations, and does not use any parallel data. This paper attempts to explicitly evaluate the crosslingual alignment and identify the key elements in transliteration-based approaches that contribute to better performance. For this, we train multiple models under varying setups for two pairs of related languages: (1) Polish and Ukrainian and (2) Hindi and Urdu. To assess alignment, we define four types of similarities based on sentence representations. Our experimental results show that adding transliterations alone improves the overall similarities, even for random sentence pairs. With the help of auxiliary transliteration-based alignment objectives, especially the contrastive objective, the model learns to distinguish matched from random pairs, leading to better crosslingual alignment. However, we also show that better alignment does not always yield better downstream performance, suggesting that further research is needed to clarify the connection between alignment and performance. The code implementation is based on https://github.com/cisnlp/Transliteration-PPA.
@@ 1 Introduction
The training of highly multilingual language models has to cope with the diversity of scripts (e.g., more than 30 in Glot500-m (ImaniGooghari et al., 2023)), which tends to reduce the effectiveness of crosslingual transfer (Dhamecha et al., 2021; Purkayastha et al., 2023). A few recent studies explore the possibility of using a common script to represent all languages (Purkayastha et al., 2023; Moosa et al., 2023) through transliteration, a process of converting the text of a language from one script to another script (Wellisch et al., 1978). The intuition is that a common script can help the model to learn more knowledge through lexical overlap since common vocabularies have been shown to contribute to better crosslinguality (Pires et al., 2019; Amrhein and Sennrich, 2020). However, this script normalization step yields models that only support one script. This also hinders efficiency, as texts have to be transliterated into the common script before being fed to the model. In addition, transliteration ambiguity (words with different meanings having the same transliteration) can be a potential problem for the effectiveness of crosslingual transfer (Liu et al., 2024c).
@@
Instead of relying solely on common-script transliterations, a recent line of work also uses transliterations as an auxiliary input to improve crosslingual alignment without expanding the vocabulary (Liu et al., 2024b; Xhelili et al., 2024). These approaches combine sentences in their original script alongside their transliteration as paired inputs for sentence- or token-level alignment objectives. Surprisingly, even without parallel data, these methods show remarkable improvement in crosslingual transfer between languages with different scripts. However, the concept of crosslingual alignment is often vaguely defined in these studies. Moreover, it remains unclear why incorporating transliterations and auxiliary alignment objectives contributes to better crosslingual alignment, which relates to the similarity between translation pairs.
@@
To this end, this work presents – to the best of our knowledge – the first attempt to explain why including transliterated data in training and incorporating transliteration-based alignment objectives, such as transliteration contrastive modeling (TCM) (Liu et al., 2024b) and transliteration language modeling (TLM) (Xhelili et al., 2024), can improve the crosslingual alignment. We first discuss definitions of crosslingual alignment (Roy et al., 2020; Hämmerl et al., 2024) and establish a clear connection between crosslingual alignment and sentence retrieval performance – an explicit evaluation of sentence-level alignment as we show in §3. We then conduct a case study on two related language pairs using different scripts: (a) Polish-Ukrainian and (b) Hindi-Urdu, to explore how and why the transliteration-augmented approach improves crosslingual alignment. Specifically, we aim to answer the following questions: (1) Does adding transliterated data alone improve crosslingual alignment? (2) How do auxiliary objectives contribute to better alignment? (3) How does alignment vary when the target language is in the original script, the Latin script, or when both the source and target languages are in the Latin script? (4) Does better alignment always lead to better downstream zero-shot crosslingual performance?
@@
To answer these questions, we define four types of similarities based on sentence-level representations and conduct a thorough analysis of how these similarities vary across multiple model variants throughout the pretraining stage. Our key experimental findings can be summarized as follows:
@@
(i) Adding transliterations alone does not improve crosslingual alignment but does enhance all types of similarities. This occurs as the similarity between randomly paired sentences also increases. However, effective alignment requires distinguishing matched pairs from random pairs. (ii) With auxiliary transliteration-based alignment objectives, transliterations serve as an intermediary: language in its original script is aligned with the transliterations of ; similarly, language in its original script is also aligned with its transliterations; transliterations of are better aligned with those of in pretraining because of increased lexical overlap. Through this process, the alignment between and (both in their original scripts) is improved. (iii) Although crosslingual alignment is generally considered crucial for enhancing zero-shot crosslingual transfer in downstream tasks, our results indicate that better alignment does not always yield better downstream performance. This finding aligns with recent findings by Hua et al. (2024) and suggests that further research is needed in the community to clarify the connection between crosslingual alignment and crosslingual transfer.
@@ 2.1 Multilingual Language Models
Models pretrained on a wide range of languages using self-supervised objectives, such as masked language modeling (MLM) (Devlin et al., 2019) or causal language modeling (Radford et al., 2019), are referred to as mPLMs. With respect to their use of the Transformer (Vaswani et al., 2017) architecture, these models can be categorized into encoder-only (Devlin et al., 2019; Conneau et al., 2020; Liang et al., 2023), encoder-decoder (Liu et al., 2020; Fan et al., 2021; Xue et al., 2021), and decoder-only models (Lin et al., 2022; Shliazhko et al., 2022; Scao et al., 2022). With the recent scale-up in both model and data size, decoder-only models, also known as large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023), can achieve impressive performance in various generation tasks across high- and medium-resource languages (Zhao et al., 2024a; Üstün et al., 2024; Zhao et al., 2024b). Parallel efforts have produced encoder-only models with very large language coverage, improving the situation for many low-resource or under-represented languages (Ogueji et al., 2021; Alabi et al., 2022; ImaniGooghari et al., 2023; Wang et al., 2023; Liu et al., 2024a). These encoder-only models excel in multiple tasks in the zero-shot crosslingual transfer manner (Huang et al., 2019; Artetxe and Schwenk, 2019; Hu et al., 2020; Zhang et al., 2024).
@@ 2.2 Training with Alignment Objectives
Most mPLMs trained without any additional crosslingual signals already show good performance across languages, possibly due to factors such as lexical overlap (Pires et al., 2019), shared position and special token embeddings (Dufter and Schütze, 2020) and even language imbalance (Schäfer et al., 2024). A recent review of factors contributing to crosslingual transfer is provided by Philippy et al. (2023). To further improve the crosslingual alignment of mPLMs, many methods additionally leverage crosslingual signals during or after pretraining. These methods can rely on bilingual dictionaries (Cao et al., 2020; Wu and Dredze, 2020; Chi et al., 2021b; Efimov et al., 2023), parallel data (Reimers and Gurevych, 2020; Pan et al., 2021; Wang et al., 2022b), or a combination of both (Wei et al., 2021; Hu et al., 2021) to facilitate crosslingual alignment. This set of methods aims to increase the similarity between paired instances (words or sentences) and sometimes also to reduce the similarity between unrelated data, via contrastive learning objectives (Chopra et al., 2005; Gao et al., 2021; Chi et al., 2021a). Another group of methods focuses on reformulating the training data, with the expectation of implicitly improving alignment through techniques such as artificial code-switching generation (Chaudhary et al., 2020; Wang et al., 2022a; Reid and Artetxe, 2022) or using a translation language modeling objective (Conneau and Lample, 2019).
@@ 2.3 Transliteration in Language Modeling
Transliteration converts the text of a language from one script to another (Wellisch et al., 1978). Since this process does not translate meanings but rather represents the original sounds as closely as possible in the target script, transliteration can be performed efficiently and accurately using a rule-based system, such as Uroman (Hermjakob et al., 2018). Recent studies have shown that better language models can be trained using data transliterated into a common script, due to improved lexical overlap (Amrhein and Sennrich, 2020; Dhamecha et al., 2021; Muller et al., 2021; Purkayastha et al., 2023; Moosa et al., 2023; Ma et al., 2024). To further break the script barrier and prevent models from supporting only one script, another line of research uses transliterations as auxiliary input to create paired data for post-pretraining with some translation-based alignment objectives (Liu et al., 2024b; Xhelili et al., 2024), resulting in better crosslingual transfer performance between languages written in different scripts. However, it remains unclear why these approaches achieve better alignment between languages written in different scripts using only transliterations as auxiliary inputs without the presence of translation data.
@@ 3 Prelimilary: Crosslingual Alignment
Definition.
Crosslingual alignment refers to the degree of similarity among representations of similar meanings across languages, which can be further classified into weak alignment and strong alignment (Roy et al., 2020).111Strong alignment requires even a greater distance of dissimilar meanings within a language, which is usually hard in multilingual NLP. Therefore it is not considered in this paper. Following the definition of weak alignment of Hämmerl et al. (2024), similar meanings (across languages) should have more similar representations than dissimilar meanings. Formally, let be a pair of representations of two units (either tokens or sentences) with similar meanings in language and respectively, weak alignment is defined as follows:
where sim is a similarity measure, such as the cosine similarity. Weak alignment requires that the representational similarity between a unit in and its (approximately) equivalent counterpart in is higher than the similarities between this unit and any other units in . It is important to note that this notion of alignment emphasizes the relative magnitude of similarity rather than the absolute magnitude. The similarity between and does not have to be very large to induce compliant alignments. Some models, though assigning similar representations to similar units, also make less related or even unrelated units similar, therefore possibly resulting in . This naturally induces suboptimal alignments, due to the failure to differentiate between similar and dissimilar meanings across languages.
Evaluations.
The definition of crosslingual alignment on the sentence level closely resembles the measure used in sentence retrieval tasks, where a model retrieves the most relevant or similar sentence in language given a query sentence in language . Therefore, sentence-level crosslingual alignment can be directly evaluated through sentence retrieval. The performance of sentence retrieval can be evaluated by calculating the top- accuracy on a given parallel corpus, using sentences from one language as the queries, and retrieving their corresponding matches in the other language.222Similarly, the word alignment task (identifying corresponding words between two texts) can be used to evaluate the crosslingual alignment, particularly at the token level. However, this task often requires high-quality, golden-labeled data, which is difficult to obtain. As a result, this study focuses solely on sentence-level crosslingual alignment. Additionally, crosslingual alignment is believed to be able to be evaluated indirectly by other downstream tasks that rely on zero-shot crosslingual transfer ability (Huang et al., 2019; Artetxe and Schwenk, 2019). That is, given an mPLM, one fine-tunes the model on the training data of a source language and then directly evaluates its performance on the test set of target languages. The underlying intuition is that models with strong alignment are often expected to perform well in such tasks, as representations of similar meanings should be consistent across languages. However, we show that better crosslingual alignment does not always lead to better downstream crosslingual performance in §5.4. This observation aligns with previous findings (Wu and Dredze, 2020; Gaschi et al., 2023), which suggest that alignment and downstream performance are not always strongly correlated.
@@ 4.1 Languages
Polish-Ukrainian pair.
Polish (pol) and Ukrainian (ukr) are Slavic languages, belonging to the West and East Slavic branch respectively. Polish and Ukrainian have historically influenced each other, contributing to shared vocabulary and linguistic features. Polish uses Latin (Latn) script while Ukrainian uses Cyrillic (Cyrl) script.
Hindi-Urdu pair.
Hindi (hin) and Urdu (urd) both belong to the Indo-Aryan branch of the Indo-European family, spoken in the Indian subcontinent. They are mostly mutually intelligible languages that historically can be viewed as two standardized dialects of Hindustani, and therefore they share large common vocabularies. A major difference is that Hindi uses the Devanagari (Deva) script while Urdu uses the Arabic (Arab) script.
An important difference between these two language pairs is that transliteration only changes the script of one language (ukr) for the pol-ukr pair, whereas it changes the script of both urd and hin for the hin-urd pair. In this way, our choices cover the most common cases, and therefore we assume the conclusions and insights from our experiments can be naturally extended to other language pairs.
@@ 4.2 Training Data
Original data.
We use the data from Glot500-c (ImaniGooghari et al., 2023) for each language of interest. For the pol-ukr pair, there are around 7M sentences for ukr_Cyrl and around 19M sentences for pol_Latn sentences. For the hin-urd pair, there are around 7M sentences for hin_Deva and 6M sentences for urd_Arab. We concatenate all data together for each language pair and refer to the final data in their original script as Data for pol-ukr and Data for hin-urd respectively.
Transliterated data.
We use Uroman (Hermjakob et al., 2018) to transliterate both Data and Data to the Latin script. We refer to the resulting Latin-script data as Data and Data for pol-ukr and hin-urd pair respectively. It is important to note that the original and transliterated data are in one-to-one correspondence. This means that the th line in Data is the transliteration of the th line in Data.
@@
4.3 Training Objectives
Masked Language Modeling (MLM).
This is the primary learning objective we use to train our model variants. This objective improves the general language modeling ability by masking certain tokens in the input sentences and learning to predict them. Following Devlin et al. (2019), we randomly replace 15% tokens in the input sentences with a special token: [mask] and use a language modeling head to reconstruct the original tokens from the final contextualized embeddings.
Transliteration Contrastive Modeling (TCM).
This contrastive objective is proposed by Liu et al. (2024b). It increases the similarity between pairs of sentence-level representations composed of one sentence in its original script and the corresponding Latin transliteration. Following Liu et al. (2024b), we obtain these representations by mean-pooling the output of the 8th Transformer layer and calculate the loss batch-wise: the positive samples are the paired sentences within a batch; the negative samples are any combinations of two sentences that are not paired within a batch.